OpenStudy (anonymous):

After two half-lives, a sample contains 10 grams of parent isotopes. How many grams of the parent isotope were present to start?

OpenStudy (kropot72):

Let x = initial number of grams of parent isotope At the end of the first half life there will be x/2 grams of parent isotope remaining. At the end of the second half life there will be \[\frac{x}{2}\times \frac{1}{2}=\frac{x}{4}\] grams of parent isotope remaining. \[\frac{x}{4}=10\] x = 40 grams of parent isotope to start with.